1. A metallic sphere of radius 4.2 cm is melted and recast into the shape of a cylinder of radius 6 cm. Find the height of the cylinder.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
It is given that radius of the sphere (R) = 4.2 cm Also, Radius of cylinder (r) = 6 cm Now, let height of cylinder = h It is given that the sphere is melted into a cylinder. So, Volume of Sphere = Volume of Cylinder ∴ (4/3)×π×R3 = π×r2×h. h = 2.74 cm